# May I Ask a Follow-up Question? Understanding the Benefits of Conversations in Neural Network Explainability

### v1

@@ 6. Conclusion
In our work, we conduct Wizard-of-Oz experiments to investigate how free-form conversations assist users in understanding static explanations, promoting trust, and making informed decisions about AI models. Participants engage in conversational explanations with XAI experts with the aim of understanding how the provided static explanation explains the model decision. To evaluate the effects of conversations, we design objective and subjective measurements. We observe a notable improvement in users’ comprehension, acceptance, trust, and collaboration after conversations. From collected conversations, we find that participants’ questions and confusions are diverse and unanticipated. Our findings advocate for the integration of dialogue systems in future XAI designs to ensure more personalized explanations. We hope that our study design and insights will aid the development of conversational XAI research.

### v2

@@ 5. Conclusion
In our work, we conduct Wizard-of-Oz experiments to investigate how free-form conversations assist users in understanding static explanations, promoting trust, and making informed decisions about AI models. Participants engage in conversational explanations with XAI experts to understand how the provided static explanation explains the model decision. To evaluate the effects of conversations, we design objective and subjective measurements. We observe a notable improvement in users’ comprehension, acceptance, trust, and collaboration after conversations. From collected conversations, we find that participants’ questions and confusions are diverse and unanticipated. Our findings advocate for the integration of dialogue systems in future XAI designs to ensure more personalized explanations.